OpenStudy (anonymous):

HELP!!!! URGENTLY!!! Find d^2y/dx^2 for y = x^4/600, at x = e a. 0.02e b. 0.02e2 c. 0.01e2 d. 0.02 e. 0.01e

OpenStudy (anonymous):

First, find y' then y''

OpenStudy (anonymous):

\[ y= \frac {x^4} {600}\\ y'= 4\frac {x^3} {600}=\frac {x^3} {150} \]

OpenStudy (anonymous):

\[ y'= 4\frac {x^3} {600}=\frac {x^3} {150}\\ y'' =3 \frac {x^2} {150}\\= \frac {x^2} {50} \]

OpenStudy (anonymous):

Replace x by e above and you are done.

OpenStudy (anonymous):

\[ \frac {e^2}{50}=\frac {2e^2}{100}= .02 e^2 \]

OpenStudy (anonymous):

Answer b

OpenStudy (anonymous):

Thank you so much! I really appreciate it! :-)